# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v7

@@ VII-B3 Sim-to-Real Paradigms
Real2Sim2real [363] enhanced imitation learning in real-world scenarios by leveraging reinforcement learning (RL) trained in a “digital twin” simulation environment. This method involves strengthening strategies through extensive RL within the simulation, followed by transferring these strategies to the real world to address data scarcity and achieve effective robotic operational imitation learning. Initially, NeRF and VR are used for scene scanning and reconstruction, and the constructed scene assets are imported into the simulator to achieve real-to-simulation fidelity. Subsequently, RL in the simulation fine-tunes the initial strategies derived from sparse expert demonstrations collected in the real world. Finally, the refined strategies are transferred to real-world settings.
TRANSIC [364] narrowed the sim-to-real gap by enabling real-time human intervention to correct robot behaviors in real-world scenes. It enhances sim-to-real transfer performance through several steps: first, robots are trained using RL to establish foundational strategies within a simulation environment. Then, these strategies are implemented on real robots, with humans intervening and correcting behaviors in real-time via remote control when errors occur. The data collected from these interventions are used to train a residual policy. Integrating both foundational and residual policies ensures smoother trajectories in real-world applications following sim-to-real transfer. This approach significantly reduces the need for real-world data collection, thereby mitigating the burden while achieving effective sim-to-real transfer.
Domain Randomization [365, 366, 367] enhanced the generalization of models trained in simulated environments to real-world scenarios by introducing parameter randomization during simulation. While both simulated and real environments involve perception through camera-acquired visual images, differences such as object friction and gloss make accurate simulation challenging. Randomizing parameters during simulation training covers a wide range of conditions, potentially encompassing variations that might occur in real-world settings. This approach boosts the robustness of trained models, enabling deployment from simulation to real environments.
System Identification [368, 369] constructed an accurate mathematical model of physical scenes in real-world environments, encompassing parameters such as dynamics and visual rendering. It aims to make simulation environments closely resemble real-world settings, facilitating smooth transitions of models trained in simulation to real environments.
Lang4sim2real [370] used natural language as a bridge to address the sim-to-real gap by using textual descriptions of images as a cross-domain unified signal. This approach aids in learning domain-invariant image representations, thereby improving generalization performance across simulation and real environments. Initially, an encoder is pretrained on image data annotated with cross-domain language descriptions. Subsequently, using the domain-invariant representations, a multi-domain, multi-task language-conditioned behavioral cloning policy is trained. This method compensates for the scarcity of real-world data by additional information from abundant simulated data, thereby enhancing sim-to-real transfer.
VII-C Embodied Control
Embodied control learns through interaction with the environment and optimizes behavior using a reward mechanism to obtain the optimal policy, thereby avoiding the drawbacks of traditional physical modeling methods. Embodied control methods can be divided into two types:
1) Deep Reinforcement Learning (DRL). DRL can handle high-dimensional data and learn complex behavior patterns, making it suitable for decision-making and control. The hybrid and dynamic policy gradient (HDPG) [371] is proposed for biped locomotion, allowing the control policy to be simultaneously optimized by multiple criteria dynamically. DeepGait [372] is a neural network policies for terrain-aware locomotion, which combines methods for model-based motion planning and reinforcement learning. It includes a terrain-aware planner for generating gait sequences and base motions guiding the robot towards target directions, along with a gait and base motion controller for executing these sequences while maintaining balance. Both the planner and controller are parameterized using neural network function approximators and optimized using deep reinforcement learning algorithms.
Figure 18: Examples of embodied control for various locomotion modes, demonstrating the robots’ agile movement and interaction capabilities. The orange box on the left showcases the locomotion of quadruped robots (Unitree B1, ANYmal, Spot, Ghost Robotics Vision-60, MIT Cheetah) and one biped robot (ATRIAS). The red box on the middle displays motion scenes of the Atlas Robot climbing stairs, TELESAR V robot writing, Unitree H1 high-fiving, and the Boston robot leaping. The green box shows multi-robot collaborative control, finding path and planning path.
Figure 19: All robots in one. The figure is referred from the work [387]. The key advantages of the ARIO standard and dataset can be summarized as follows: 1) Comprehensive Sensory Modalities: ARIO incorporates five sensory modalities—images, 3D data, sound, text, and tactile information—offering a richer and more diverse dataset, 2) Timestamp-Based Data Alignment: Data is recorded and synchronized based on timestamps, accommodating different sensor frame rates, 3) Structured Series-Task-Episode Framework: ARIO uses a clear series-task-episode structure, with detailed textual descriptions for each series and task, 4) Unified Data Format: The dataset follows a standardized format, supporting various robot types and control objects, simplifying data processing, 5) Integration of Simulation and Real Robot Data: ARIO includes both simulation and real-world data across multiple tasks, scenarios, and robot types, enhancing generalization across different hardware platforms.
2) Imitation Learning. The DRL has its drawbacks of requiring a large amount of data from numerous trials. To address this issue, imitation learning was introduced, which aims to minimize data usage by collecting high-quality demonstrations. To improve data efficiency, Offline RL + Online RL was proposed to reduce interaction costs and ensure safety. This method first employs offline RL to learn policies from static, pre-collected large datasets. These policies are then deployed in the real environment for real-time interaction and exploration, with adjustments made based on feedback. The representative imitation learning methods from human demonstrations are ALOHA [373] and Mobile ALOHA [359].
Although embodied AI encompasses high-level algorithms, models, and planning modules, its most fundamental and essential component is embodied control. Therefore, it is imperative to consider how to control physical entities and endow them with physical intelligence. Embodied control is closely related to hardware, such as controlling joint movements, end-effector positions, and walking speeds. For robotic arms, knowing the end-effector’s position, how to plan joint trajectories to move the arm to the target? For humanoid robots, knowing the motion patterns, how to control the joints to achieve the target posture? These are critical issues that need to be addressed in control. Several works focus on robotic control, enhancing the flexibility of robotic actions. [374] proposed a vision-based full-body control framework. By connecting a robotic arm and a robotic dog, utilizing all degrees of freedom (12 joints in legs, 6 joints in the arm, and 1 in the gripper), it tracks the robot dog’s speed and the robotic arm’s end-effector position, achieving more flexible control. Some works [375, 376] employed traditional methods to control bipedal robot walking. MIT’s Cheetah 3 [377], ANYmal [378], and Atlas [379] used robust walking controllers to manage the robots. These robots can be used for more agile motion tasks, such as jumping or overcoming various obstacles [380, 381, 382, 383, 384]. Other works [385, 386] focused on the control of humanoid robots, to perform various actions like humans and mimic human behaviors. Fig .18 illustrates some examples.
Embodied control integrates RL and sim-to-real techniques to optimize strategies through environmental interaction, enabling the exploration of unknown domains, potentially surpassing human capabilities, and adapting to unstructured environments. While robots can mimic many human behaviors, effective task completion often requires RL training based on environmental feedback. The most challenging scenarios include contact-intensive tasks, where manipulation necessitates real-time adjustments based on feedback such as the state, deformation, material, and force of manipulated objects. In such cases, RL is indispensable. In the era of MLMs, these models possess a generalized understanding of scene semantics, providing robust reward functions for RL. Additionally, RL is crucial for aligning large models with their intended tasks. In the future, after pre-training and fine-tuning, RL is still required to align with the physical world, ensuring effective deployment in real-world environments.
VII-D All Robots in One
@@
Significant data-level limitations continue to impede the development of robust, general-purpose embodied agents, particularly concerning standardized formats, diversity, and data volume. Task-specific datasets are insufficient for training these versatile agents. Despite the seemingly unified structure of pre-training datasets like Open X-embodiment, several critical issues remain unresolved. These issues include the absence of comprehensive sensory modalities—no dataset currently integrates images, 3D vision, text, tactile, and auditory inputs simultaneously. Additionally, the lack of a unified format in multi-robot datasets complicates data processing and loading. Furthermore, there is an incompatibility in representing diverse control objects across different robotic platforms, insufficient data volume that hinders large-scale pretraining, and a scarcity of datasets that combine both simulated and real data, which is essential for addressing the sim-to-real gap.
@@
To overcome these challenges, ARIO (All Robots In One) [387] is introduced, which is a new dataset standard that optimizes existing datasets and facilitates the development of more versatile and general-purpose embodied AI agents. The ARIO standard records control and motion data from robots with different morphologies in a unified format. A timestamp mechanism standardizes data collection by addressing variations in robot action frequencies and sensor frame rates. ARIO’s unified format accommodates variable data from diverse robot types, ensuring precise timestamps. This standard enables users to efficiently train high-performing, generalizable embodied AI models, positioning ARIO as the ideal format for embodied AI datasets.
Building on the ARIO standard, a unified large-scale ARIO dataset is further developed, which comprises approximately 3 million episodes collected from 258 series and 321,064 tasks. This dataset was assembled through a multi-faceted approach: (1) real-world data collection from the custom platform, yielding 3,662 episodes across more than 105 tasks, (2) simulation-based data generation using platforms such as Habitat, MuJoCo, and SeaWave, resulting in 703,088 episodes from 1,198 tasks, and (3) the conversion of existing open-source datasets to the ARIO standard, contributing 2,326,438 episodes from 319,761 tasks. The ARIO dataset addresses the limitations of current datasets and facilitates the development of robust, general-purpose embodied agents. By providing a cohesive framework for data collection and representation, ARIO paves the way for the development of more powerful and versatile embodied AI agents, capable of navigating and interacting with the physical world in increasingly complex and diverse ways. The project is available on https://imaei.github.io/project_pages/ario.
@@ VIII Challenges and Future Directions
High-quality Robotic Datasets: Obtaining sufficient real world robotic data remains a significant challenge. Collecting this data is both time-consuming and resource-intensive. Relying solely on simulation data worsens the sim-to-real gap problem. Creating diverse real world robotic datasets necessitates close and extensive collaboration among various institutions. Additionally, the development of more realistic and efficient simulators is essential for improving the quality of simulated data. Current work RT-1 [11] used pre-trained models based on robot images and natural language commands. RT-1 has achieved good results in navigation and grasping tasks, but acquiring real world robot datasets is very challenging. For building generalizable embodied models capable of cross-scenario and cross-task applications in robotics, it is essential to construct large-scale datasets, leveraging high-quality simulated environment data to assist real world data.
Efficient Utilization of Human Demonstration Data: Efficient utilization of human demonstration data involves leveraging the actions and behaviors demonstrated by humans to train and improve robotic systems. This process includes collecting, processing, and learning from large-scale, high-quality datasets where humans perform tasks that robots are intended to learn. Current work R3M [388] used action labels and human demonstration data to learn generalizable representations with high success rates, but the efficiency for complex tasks still needs improvement. Therefore, it is important to effectively utilize large amounts of unstructured, multi-label, and multi-modal human demonstration data combined with action label data to train embodied models that can learn various tasks in relatively short periods. By efficiently utilizing human demonstration data, robotic systems can achieve higher levels of performance and adaptability, making them more capable of performing complex tasks in dynamic environments.
Cognition of Complex Environment: Cognition of complex environment refers to the ability of embodied agents in physical or virtual environments, to perceive, understand, and navigate complex real world environments. Based on extensive commonsense knowledge, the Say-Can [299] utilized pre-trained LLM models’ task decomposition mechanism, which relies heavily on large amounts of commonsense knowledge for simple task planning but lacking understanding of long-term tasks in complex environments. For unstructured open environments, current works usually rely on pre-trained LLMs’ task decomposition mechanism using extensive common-sense knowledge for simple task planning, while lacking specific scene understanding. It is vital to enhance the ability of knowledge transfer and generalization in complex environments. A truly versatile robotics system should be capable of comprehending and executing natural language instructions across diverse and unseen scenes. This necessitates the development of adaptable and scalable embodied agent architectures.
@@
Long-Horizon Task Execution: Executing single instructions can often entail long-horizon tasks for robots, exemplified by commands like “clean the kitchen”, which involve activities such as rearranging objects, sweeping floors, wiping tables, and more. Accomplishing such tasks successfully necessitates the robot’s ability to plan and execute a sequence of low-level actions over extended time spans. While current high-level task planners have shown initial success, they often prove inadequate in diverse scenarios due to their lack of tuning for embodied tasks. Addressing this challenge requires the development of efficient planners equipped with robust perception capabilities and much commonsense knowledge.
@@
Causal Relation Discovery: Existing data-driven embodied agents make decisions based on the intrinsic correlations within the data. However, this modeling approach does not allow the models to truly understand the causal relations between knowledge, behavior, and environment, resulting in biased strategies. This makes it difficult to ensure that they can operate in real-world environments in an interpretable, robust, and reliable manner. Therefore, it is important for embodied agents to be driven by world knowledge, capable of autonomous causal reasoning. By understanding the world through interaction and learning its workings via abductive reasoning, we can further enhance the adaptability, decision reliability, and generalization capabilities of multimodal embodied agents in complex real-world environments. For embodied tasks, it is necessary to establish spatial-temporal causal relations across modalities through interactive instructions and state predictions [389]. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [390], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied AI.
Continual Learning: In robotics applications, continual learning [391] is crucial for deploying robot learning policies in diverse environments, yet it remains a largely unexplored domain. While some recent studies have examined sub-topics of continual learning—such as incremental learning, rapid motor adaptation, and human-in-the-loop learning—these solutions are often designed for a single task or platform and do not yet consider foundational models. Open research problems and viable approaches include: 1) mixing different proportions of prior data distribution when fine-tuning on the latest data to alleviate catastrophic forgetting [392], 2) developing efficient prototypes from prior distributions or curricula for task inference in learning new tasks, 3) improving training stability and sample efficiency of online learning algorithms, 4) identifying principled ways to seamlessly incorporate large-capacity models into control frameworks, potentially through hierarchical learning or slow-fast control, for real-time inference.
@@
Unified Evaluation Benchmark: While numerous benchmarks exist for evaluating low-level control policies, they often vary significantly in the skills they assess. Furthermore, the objects and scenes included in these benchmarks are typically limited by simulator constraints. To comprehensively evaluate embodied models, there is a need for benchmarks that encompass a diverse range of skills using realistic simulators. Regarding high-level task planners, many benchmarks focus on assessing planning capability through question-answering tasks. However, a more desirable approach involves evaluating both the high-level task planner and the low-level control policy together for executing long-horizon tasks and measuring success rates, rather than relying solely on isolated assessments of the planner. This integrated approach offers a more holistic assessment of the capabilities of embodied AI systems.
@@ IX Conclusion
Embodied AI allows agents to sense, perceive, and interact with various objects from both cyber space and physical world, which exhibits its vital significance toward achieving AGI. This survey extensively reviews embodied robots, simulators, four representative embodied tasks: visual active perception, embodied interaction, embodied agents and sim-to-real robotic control, and future research directions. The comparative summary of the embodied robots, simulators, datasets, and approaches provides a clear picture of the recent development in embodied AI, which greatly benefits the future research along this emerging and promising research direction.
Acknowledgement
We sincerely thank Jingzhou Luo, Xinshuai Song, Kaixuan Jiang, Zhida Li, and Ganlong Zhao for their contributions.

### v8

@@ VII-B3 Sim-to-Real Paradigms
Real2Sim2real [229] improves imitation learning by using reinforcement learning in a ”digital twin” simulation to develop strategies, which are then transferred to the real world. TRANSIC [230] reduces the sim-to-real gap through real-time human intervention and residual policy training based on corrected behaviors. Domain Randomization [231, 232, 233] increases model generalization by varying simulation parameters to cover real-world conditions. System Identification [234, 235] creates accurate simulations of real-world scenes to ensure smooth transitions from simulation to reality. Lang4sim2real [236] leverages natural language descriptions to bridge the sim-to-real gap, improving model generalization with cross-domain image representations.
Figure 12: Exemplar tasks from ARIO, where the top row indicates the task category while the text at the bottom row provides task instructions.
VII-B4 ARIO (All Robots in One)
@@
Despite the seemingly unified structure of pre-training datasets like Open X-embodiment, several critical issues remain unresolved. These issues include the absence of comprehensive sensory modalities—no dataset currently integrates images, 3D vision, text, tactile, and auditory inputs simultaneously. Additionally, the lack of a unified format in multi-robot datasets complicates data processing and loading. Furthermore, there is an incompatibility in representing diverse control objects across different robotic platforms, insufficient data volume that hinders large-scale pretraining, and a scarcity of datasets that combine both simulated and real data, which is essential for addressing the sim-to-real gap.
@@
To overcome these challenges, we propose ARIO (All Robots In One) , which is a new dataset standard that optimizes existing datasets and facilitates the development of more versatile and general-purpose embodied agents. The ARIO standard [237] records control and motion data from robots with different morphologies in a unified format. ARIO’s unified format accommodates variable data from diverse robot types, ensuring precise timestamps. This standard enables users to efficiently train high-performing, generalizable embodied AI models, positioning ARIO as the ideal format for embodied AI datasets. Building on the ARIO standard, a unified large-scale ARIO dataset is further developed, which comprises approximately 3 million episodes collected from 258 series and 321,064 tasks. Fig. 12 shows exemplar tasks from ARIO.
The ARIO dataset addresses the limitations of current datasets and facilitates the development of robust, general-purpose embodied agents. By providing a cohesive framework for data collection and representation, ARIO paves the way for the development of more powerful and versatile embodied agents, capable of navigating and interacting with the physical world in complex and diverse ways.
VII-B5 Real-world Deployments of Embodied AI Systems
Embodied AI systems have made significant strides across various scenes. In healthcare, robots like the Da Vinci Surgical System and Moxi automate tasks such as surgery precision and supply deliveries. In logistics, Amazon Robotics and Boston Dynamics’ Stretch improve efficiency in warehousing and transportation. Manufacturing benefits from AI-driven robots like Fanuc and ABB, enhancing precision and collaboration. Nevertheless, sim-to-real adaptation faces significant challenges, including the domain gap between simulation and real-world data distributions, the complexity of dynamic real-world interactions, and the limited diversity of training data. Models often overfit to simulations, struggle with real-world sensor noise, and fail to handle unexpected events.
@@ VIII Challenges and Future Directions
High-quality Robotic Datasets: Obtaining sufficient real world robotic data remains a significant challenge. Collecting this data is both time-consuming and resource-intensive. Relying solely on simulation data worsens the sim-to-real gap problem. Creating diverse real world robotic datasets necessitates close and extensive collaboration among various institutions. Additionally, the development of more realistic and efficient simulators is essential for improving the quality of simulated data. For building generalizable embodied models capable of cross-scenario and cross-task applications in robotics, it is essential to construct large-scale datasets, leveraging high-quality simulated environment data to assist real world data.
@@
Long-Horizon Task Execution: Executing single instructions can often entail long-horizon tasks for robots, exemplified by commands like “clean the kitchen”, which involve activities such as rearranging objects, sweeping floors, wiping tables, and more. Accomplishing such tasks successfully necessitates the robot’s ability to plan and execute a sequence of low-level actions over extended time spans. While current high-level task planners have shown initial success, they often prove inadequate in diverse scenarios due to their lack of tuning for embodied tasks. Addressing this challenge requires the development of efficient planners equipped with robust perception capabilities and much commonsense knowledge. To balance the trade-off between planning complexity and real-time adaptability, we can combine lightweight monitor modules for high-frequency monitoring, and two adapters for subtask and path adaptation reasoning at a lower frequency.
@@
Causal Reasoning: Existing data-driven embodied agents make decisions based on intrinsic data correlations. However, this approach does not allow the models to truly understand the causal relations between knowledge, behavior, and environment, resulting in biased strategies. This makes it difficult to ensure that they can operate in real-world environments in a robust and reliable manner. Therefore, it is important for embodied agents to be driven by world knowledge, capable of autonomous causal reasoning. By understanding the world through interaction and learning its workings via abductive reasoning, we can further enhance the adaptability, decision reliability, and generalization capabilities of embodied agents in complex real-world environments. For embodied tasks, it is necessary to establish spatial-temporal causal relations across modalities through interactive instructions and state predictions. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning in dynamic scenes.
@@
Unified Evaluation Benchmark: While numerous benchmarks exist for evaluating low-level control policies, they vary significantly in the assessed skills. Furthermore, the objects and scenes in these benchmarks are typically limited by simulator constraints. To comprehensively evaluate embodied models, the benchmark should encompass a diverse range of skills using realistic simulators. Many benchmarks for high-level task planners focus on assessing planning capability through question-answering tasks. However, a more desirable approach involves evaluating both the high-level task planner and the low-level control policy together for executing long-horizon tasks and measuring success rates, rather than relying solely on isolated assessments of the planner.
Security and Privacy: Embodied agents face significant security challenges when deployed in sensitive or private spaces. These agents often rely on LLMs for decision-making, which introduces new vulnerabilities. For instance, LLMs are susceptible to backdoor attacks like word injection, scenario manipulation, and knowledge injection, which can lead to dangerous outcomes such as autonomous vehicles accelerating towards obstacles or robots performing hazardous actions. To mitigate these risks, we can evaluate potential attack vectors and develop more robust defenses. Additionally, the secure prompting, state management, and safety validation mechanisms can be used to enhance the security and robustness.
@@ IX Conclusion
Embodied AI allows agents to sense, perceive, and interact with various objects from both cyber space and physical world, which exhibits its vital significance toward achieving AGI. This survey extensively reviews embodied robots, simulators, four representative embodied tasks: visual active perception, embodied interaction, embodied agents and sim-to-real adaptation, and future research directions. The comparative summary of the embodied robots, simulators, datasets, and approaches provides a clear picture of the recent development in embodied AI, which greatly benefits the future research along this emerging and promising research direction.